Normal forces are applied uniformly over the surface of a spherical volume of ethanol whose radius is 12.0 cm. If the pressure on the surface is increased by 150 MPa, by how much (in cm) does the radius of the sphere decrease?
Answer #1

Volume () R AV V-Vo 47T Ro B=-AP Vo AV AP V 3 4R2 R A(R 41T - R Ro AR= R-Ro Rol- Ro 47 -(acm) AR- Ro (1- 150xISPa 022xi0a - L